OpenStudy (anonymous):

Solve the following system of equations. 4x+5y = 5 8x+5y = 25 i need to find what x and y equals

pooja195 (pooja195):

You need to find a common factor first 4x+5y = 5 8x+5y = 25 In this case it is -2 :) Bascially ask yourself which number will i need to multiply to get opposite x values that will cancel out \(\huge\rm\color{red}{-2(4x+5y = 5)}\) Tell me the equation you get once you distribute

OpenStudy (anonymous):

-8x-10y=-10 not sure if this is right I'm kinda bad at math @pooja195

pooja195 (pooja195):

Thats correct good! :D \[\huge~\rm-8x-10y=-10\] \[\huge~\rm~~~~~8x+5y = 25 \] Now just do simple math whats \(\huge\rm\color{red}{-8+8=?}\)

OpenStudy (anonymous):

-8 +8 is 0 lol

pooja195 (pooja195):

Good so that cancels out :) -10+5= ?

pooja195 (pooja195):

Need help?

OpenStudy (anonymous):

-10+5 = -5???

pooja195 (pooja195):

good so we have -5y=15 i added the -10+25 ok now that we have that -5y=15 we divde both sides by -5 -5 div -5=1 therefore x is alone now :P whats 15/-5 ?

OpenStudy (anonymous):

15/-5 = -3?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Yes thereofore our y value is -3 now we use substituion to figure out our x value \[\huge~\rm −8x−10(-3)=−10\] solve for x

OpenStudy (anonymous):

im kinda stuck to be honest -.- I'm so bad at math

pooja195 (pooja195):

\[\huge~\rm~−8x−(10)(−3)=−10\] Simplify both sides of the equation. \[\huge~\rm~−8x+30=−10\] Subtract 30 from both sides. \[\huge~\rm~−8x+30−30=−10−30\]\[\huge~\rm~−8x=−40\] Divide both sides by -8 Take it from here :)

OpenStudy (anonymous):

so x would be 5 right ?

pooja195 (pooja195):

yes! :D good job :) (5,-3)

OpenStudy (anonymous):

thank you so much

pooja195 (pooja195):

No problem :D